The probabilities of a machine manufacturing 0,1,2,3,4 or 5 defective parts in one day are 0.75,0.17,0.04,0.025,0.01, and 0.005, respectively. Find the mean, variance and standard deviation of the probability distribution. Round your answer to the nearest hundredths. 


1
Expert's answer
2021-03-01T07:11:21-0500

We know that,

"\text{Mean}(\mu) =\Sigma_{i=0} ^{5} x_iP(x_i)"

"Variance(\sigma^2)= \Sigma_{i=0}^5 x^2_{i}P(x_i)-\mu^2"

"Standard Deviation= \sqrt{\sigma^2}"




"\text{Mean}(\mu) =\Sigma_{i=0} ^{5} x_iP(x_i)= 0.39"


"Variance(\sigma^2)= \Sigma_{i=0}^5 x^2_{i}P(x_i)-\mu^2\\"

"\Rightarrow0.84-0.1521=0.6879 = 0.69"


"Standard Deviation= \sqrt{\sigma^2}"

"\Rightarrow 0.829=0.83"